OpenStudy (anonymous):

this is an equation of a horizontal hyberpola, true or false (x-2)^2/36-(y+1)^2/64=1

OpenStudy (anonymous):

True

OpenStudy (anonymous):

@rini94 You can always use WolframAlpha to help with graphing.

OpenStudy (anonymous):

It's horizontal hyperbola, because x is positive!

OpenStudy (anonymous):

the focus for this parabola is (-2,0) x=-1/8y^2 true or false

OpenStudy (anonymous):

I'm back!

OpenStudy (anonymous):

Be aware that x = 4p y^2 is horizontal parabola!

OpenStudy (anonymous):

4p = -1/8 => p = -1/32 Thus focus ( h + p, k) = (-1/32, 0 )